OpenStudy (anonymous):

Which value of x makes the following equation true? 2(x + 6) = x + 16 10 16 -2 4

OpenStudy (the_fizicx99):

Distribute first, 2x + 12 = x +16 combine like terms. 2x + 12 = x +16 -x + 12 = -x + 16 x +12 = 16 -12 -12 x = 4

OpenStudy (the_fizicx99):

Geddit? :o

OpenStudy (anonymous):

omg thanks so freaken much :D

OpenStudy (the_fizicx99):

My pleasure!